Sol Summary – February 7th

Crew 188 Sol 10 Summary Report 7FEB2018

Sol 10

Summary Title: Game On!
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: Crew feeling good after a full day.

Sol Activity Summary:
Even with one EVA, the crew had a very busy and productive day. The EVA challenged the crew to a new slew of Martian Olympiad events counting down to the Winter Olympic Games back on Earth. An opening ceremony occurred and the crew hopes to share this artistic footage in high resolution 360 views. There are sure to be phots posted across social media, just look for #MDRS188. Meanwhile in the Hab, the NASA Glenn dust load experiment continued and air filtration dust samples were obtained from the EVA crew returning into the airlock. The afternoon was highlighted with staring at the Sun with a Hydrogen-alpha filter seeing sunspots, flares and even an airplane captured on camera passing in front of the Sun! The crew was busy digitally capturing video footage and audio for podcasts, school groups and for a mission documentary. After sunset the crew enjoyed watching the International Space Station pass-by briefly and an Iridium satellite. Who knows what space infrastructure Mars inhabitants will be able to watch from their homes, but they will need some form of planetary connectivity.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
Two EVAs are panned that will fill most of the operational day. Afterwards the third and final human factors group chat will occur.

Anomalies in work:
Propane was detected in the afternoon. The crew is carefully monitoring and performing basic bubble testing on pipes. Currently no leaks are detected.

Calm winds and clear skies with temperatures dropping rapidly at sunset.

Crew Physical Status:
One crewmember has cold symptoms, which may be due to fatigue but health status is nominal.

Two are planned for tomorrow.

Reports to file:
1. Ops Report
2. Sol Summary
3. Journalist Report (combined Sol 9 and 10)
4. Green Hab Report
5. EVA #13 and #14 request
6. EVA #12 Reports
7. Daily Photos

Support Requested (same as previous report):
– Heater for science dome requested. A purchase order request was made.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open. [end]

Copyright © The Mars Society. All rights reserved. | Main Site